OpenStudy (josee):

Find the x-intercept of the tangent line to the graph of f at the point P(0,f(0)) when f(x)= e^-x(5cosx+2sinx)

OpenStudy (anonymous):

find \[f'(x)\] then find \[f'(0)\] for the slope of the line. the equation will then be \[y=f'(0)x+f(0)\]

OpenStudy (anonymous):

\[f(0)=e^0(5\cos(0)+2\sin(0))=5\]

OpenStudy (anonymous):

\[f'(x)=-e^{-x}(5\cos(x)+2\sin(x))+e^{-x}(-5\sin(x)+2\cos(x))\] by the product rule so \[f'(0)=-5+2=-3\] if my arithmetic is correct

OpenStudy (anonymous):

and therefore the equation for the line is \[y=-3x+5\] and so the x- intercept is \[0=-3x+5\] \[x=\frac{5}{3}\]